OpenStudy (anonymous):

Which equation is a quadratic equation? Answer y + 3 = 5x + 10 y + 4x^2 = (4x - 2)(x + 3) y - 8x = (x^2 + 1)(9x - 7) y = 2x + 7x(x + 1)

sam (.sam.):

expand all and see whether which one has the highest degree of x^2

OpenStudy (anonymous):

isnt there a simpler way to see if the equation is quadratic?

sam (.sam.):

yes

sam (.sam.):

you realize that in y = 2x + 7x(x + 1) , you expand it you'll get x^2

OpenStudy (anonymous):

yea

sam (.sam.):

others like y - 8x = (x^2 + 1)(9x - 7) y + 3 = 5x + 10 y + 4x^2 = (4x - 2)(x + 3) will give you x^3 or x^4 so that's incorrect

sam (.sam.):

you want quadratic, so, y = 2x + 7x(x + 1) expanded you'll get a quadratic

OpenStudy (anonymous):

k thanks